Payback Period, IRR, and Minimum Cash Flows
The management of Mesquite Limited is currently evaluating the following investment proposal:

Time 0 Year 1 Year 2 Year 3 Year 4
Initial investment $270,000 -- -- -- --
Net operating
cash inflows -- $100,000 $100,000 $100,000 $100,000

(a) Determine the proposal's payback period.
_____ years (Round answer to one decimal place.)

(b) Determine the proposal's internal rate of return. (Refer to Appendix 12B if you use the table approach.)
Round answer to the nearest percentage (ex: 0.18567 = 19%)

     _____ %

(c) Given the amount of the initial investment, determine the minimum annual net cash inflows required to obtain an internal rate of return of 8 percent. Round the answer to the nearest dollar.

   $ ______
Answer #1

a. Payback period = Initial investment / Annual cash inflow

= $270000/100000

= 2.7 Years

b. Cumulative present value factor = 270000/100000 = 2.700

IRR = 18%

c. Cumulative present value factor at 8% for 4 years = 3.312

Minimum annual net cash inflows = 270000/3.312 = $81,522